40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 23:59:22 2020, anchal, Summary, ALS, ALS noise budget update ALS_nb_ExtraPoles.pdfALS_controls.yamlALS_nb_Kiwamus_Values.pdfALS_controls_Kiwamus_Values.yamlALS_simulink_model.svg
    Reply  Sun Sep 20 23:12:13 2020, rana, Summary, ALS, ALS noise budget update 
       Reply  Tue Sep 22 00:14:43 2020, anchal, Summary, ALS, ALS noise budget update ALS_NoiseBudgetUpdate.pdfALS_controls.yaml
          Reply  Tue Sep 22 12:14:42 2020, rana, Summary, ALS, ALS noise budget update 
             Reply  Wed Sep 23 11:13:49 2020, anchal, Summary, ALS, ALS noise budget update ALS_NoiseBudgetUpdate.pdfALS_controls.yaml
                Reply  Wed Oct 7 16:56:23 2020, anchal, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values Extract_X_AUX_PDH_Model.zipALS_NoiseBudgetUpdate.pdf
                   Reply  Thu Oct 8 11:59:52 2020, rana, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values 
                      Reply  Fri Oct 9 18:32:14 2020, anchal, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values 
                         Reply  Wed Oct 14 17:03:55 2020, anchal, Summary, ALS, ALS noise budget update - Added whitening filter for ADC ALS_NoiseBudgetUpdate.pdf
Message ID: 15617     Entry time: Wed Oct 7 16:56:23 2020     In reply to: 15601     Reply to this: 15619
Author: anchal 
Type: Summary 
Category: ALS 
Subject: ALS noise budget update - Updated AUX PDH Loop values 

AUX PDH Loop update

I used D1400293 to get the latest logged details about the universal PDH box used to lock the green laser at X end. The uPDH_X_boost.fil file present there was used to obtain the control model for this box. See attachment one for the code used. Since there is a variable gain stage in the box, I tuned the gain of the filter model F_AUX in ALS_controls.yml to get the maximum phase margin in the PDH lock of the green laser. Unity gain frequency of 8.3 kHz can be achieved in this loop and as Gautam pointed out earlier, it can't be increased much further without changes in the box.

ALS Noise Budget update

The ALS control model remains stable with a reduction in total estimate noise because of the above update. There are few things to change though:

  • This model is for a single arm locking where the beatnote signal between green laser and frequency doubled main laser is fed back to ETM at X end. Currently, Gautam is using a different scheme to lock where the feedback is sent to PSL-MC loop and the beat is taken between IR signals.
  • In the LSC controls, I couldn't find a place where the digital ALS filter I have been optimizing and Kiwamu used, was placed. From what I gathered, after demodulation of beat note signal, a digital PLL is employed and the error signal is few to the Servo Filters directly. I might be missing some script which specifically switches on a particular set of filter modules in the XARM/YARM path when arms are locked through ALS.
  • Another straight forward job for me is to verify the PSL-MC loop parameters with he TTFSS used. I'll do this next.
Attachment 1: Extract_X_AUX_PDH_Model.zip  2 kB  Uploaded Wed Oct 7 18:03:04 2020
Attachment 2: ALS_NoiseBudgetUpdate.pdf  192 kB  Uploaded Wed Oct 7 18:03:31 2020  | Hide | Hide all
ALS_NoiseBudgetUpdate.pdf ALS_NoiseBudgetUpdate.pdf ALS_NoiseBudgetUpdate.pdf ALS_NoiseBudgetUpdate.pdf ALS_NoiseBudgetUpdate.pdf
ELOG V3.1.3-